### v3

@@ Abstract
We describe the design, implementation, and evaluation of a low-cost (approximately \$15 per person) and scalable program, called Challenges, aimed at aiding women in Poland transition to technology-sector jobs. This program helps participants develop portfolios demonstrating job-relevant competencies. We conduct two independent evaluations, one of the Challenges program and the other of a traditional mentoring program – Mentoring – where experienced tech professionals work individually with mentees to support them in their job search. Exploiting the fact that both programs were oversubscribed, we randomized admissions and measured their impact on the probability of finding a job in the technology sector. We estimate that Mentoring increases the probability of finding a technology job within four months from 29% to 42% and Challenges from 20% to 29%, and the treatment effects do not attenuate over 12 months. Since both programs are capacity constrained in practice (only 28% of applicants can be accommodated), we evaluate the effectiveness of several alternative prioritization rules based on applicant characteristics. We find that a policy that selects applicants based on their predicted treatment effects increases the average treatment effect across the two programs to 22 percentage points. We further analyze how alternative prioritization rules compare to the selection that mentors used. We find that mentors selected applicants who were more likely to get a tech job even without participating in the program, and the treatment effect for applicants with similar characteristics to those selected by mentors is about half of the effect attainable when participants are prioritized optimally.
@@ 1 Introduction
An important concern of labor policymakers is identifying and implementing effective programs to support workers transitioning into occupations and industries that offer attractive jobs. One sector that has experienced rapid growth in recent decades and offers many attractive career opportunities is the technology sector. However, not all societal groups have benefited equally from opportunities in this sector; in particular, women remain underrepresented. Despite a number of programs and initiatives aimed at supporting women in the technology sector, [^1] For example, the United Nations organizes an annual event Girls in ICT Day; The European Commission carries out an annual study: Women in Digital Scoreboard. There are various NGOs dedicated to increasing women’s participation in the technology sector; For instance, Women in Technology International is a community of 160,000 women in the technology sector; Girls Who Code operates a network of 10,000 clubs providing support for women in technology. Finally, many major technology companies have initiatives aimed at supporting women in technology, including Women in Tech from Salesforce, The Women at Microsoft, or Accenture’s Women in TS&A. women held only 30% of technology jobs worldwide in 2021 (United Nations, 2021). The shortage of empirical evidence evaluating the effectiveness of specific strategies to address this issue is one of the hurdles in designing a successful policy response.
In this paper, we ask whether interventions can have an impact on this problem, and if so, which workers benefit most. We conduct randomized evaluations of two distinct programs, both implemented in Poland. The first, Mentoring, involves one-to-one mentoring and is thus not easy to scale. We designed and introduced a new program, Challenges, to be scalable. In Challenges, women work in teams to create portfolios to show to employers. Second, we ask which women benefit most from each program. Using data from the two evaluations, we identify the distinct subgroups with the largest benefits. Third, we evaluate the effectiveness of alternative prioritization rules for admission to the program in the presence of capacity constraints, illustrating that there are large gains from the optimal prioritization. Fourth, we examine the implicit selection rule used by mentors, who each selected two candidates from a pool of applicants. We find that mentors do not select the candidates with highest estimated treatment effects, but rather focus more on candidates who would be likely to succeed even without the program.
The results from our randomized evaluation suggest that the Challenges program increases the likelihood of successfully transitioning to the technology sector by approximately 45% within four months of graduation. [^2] Throughout this paper, we consider technology jobs as all jobs in firms in the technology sector other than those in finance, regulatory, legal, accounting, or human resources or jobs outside of the technology sector that involve software development and testing, IT support, and data analytics. Specifically, all jobs in UX and front-end programming are considered technology jobs. Challenges helps participants develop portfolios demonstrating capability relevant to the jobs they seek. The program was designed in partnership with technology companies, delivered online, and relies on interactions between program participants, which makes it cheap to operate (approximately \$15 per person) and suitable for scaling up. The program is also transferable across domains of the technology sector and languages; the first edition, evaluated in this paper, had specializations in user experience design and front-end programming. The program’s second edition was offered in Ukrainian instead of in Polish and offered additional specializations in software testing, cloud computing, and user interface design.
We designed the Challenges program as a scalable alternative to traditional in-person mentoring. To develop it, we partnered with Dare IT, a Polish organization offering a non-profit mentoring program – Mentoring, since 2018. Dare IT’s mentoring program serves women who, despite having invested time into acquiring IT skills, still have not successfully obtained a technology job. There is a high demand for this program, as evidenced by the fact that it has been repeatedly oversubscribed. However, scaling it up to meet the demand faces several challenges. First, it is difficult to find enough mentors. Mentors operate only on a volunteer basis and tend to be women in mid-seniority technology jobs. Since few women are in these positions, recruiting more mentors is difficult. Second, the one-to-one format of Mentoring makes it operationally demanding. Ensuring satisfaction, remedying conflicts, and tracking everyone are complicated tasks that require regular supervision by Dare IT staff.
The development of Challenges proceeded in two steps. First, together with Dare IT, we set up a series of interviews with HR experts and hiring managers of Polish technology companies to gain their perspectives on ways to support women candidates. Each of our interviewees mentioned the shortage of practical experience or clear signals of practical skills in women applicants’ profiles. Second, along with Dare IT and industry partners, we used the insights from the interviews to create Challenges. The program’s core consists of assignments sent every second week that, if completed, result in a job-relevant skill signal that can be added to participants’ resumes or mentioned during an interview. The program is designed based on three principles: (i.) it is delivered online, which lowers the costs; (ii.) it primarily involves interactions between participants, addressing the shortage of mentors; and (iii.) it is closely based on recent business cases. To operationalize the last principle, industry partners were recruited to help design the assignments and provide domain experts to review and critique selected submissions. We launched the registration to the program in December 2021 with 300 available spots; the program became oversubscribed within four hours.
We conducted two concurrent randomized experiments to evaluate the effectiveness of Challenges and Mentoring. Both experiments exploit the fact that the two programs were oversubscribed, so we were able to randomize access and evaluate the impact of participation in a program on the probability of having a technology job. To evaluate Mentoring, we asked mentors to select two mentee candidates (instead of the usual one), and within these pairs, we randomized into treatment and control. In the case of Challenges, we randomly assigned access to 200 spots in the program. Four months after the end of each program, we collected data on the outcomes. We estimate that participation in Mentoring increased the probability of having a job in technology from 29% to 42%, while the Challenges program increased this probability from 20% to 28%. [^3] The difference in the outcomes of the two control groups is likely because, in Mentoring, mentors selected applicants which were later randomized into the control group. The control group of Challenges is a random draw from the pool of applicants. We document that the outcomes of non-selected applicants are substantially lower than those of the applicants selected by mentors and later randomized into the control group. We repeated the evaluation 12 months after the end of the two programs and found that their impact did not attenuate over time.
We find that the two programs lead to different patterns of treatment effect heterogeneity. Mentoring is particularly beneficial for women from smaller towns that do not have a STEM degree, are mothers, and have been active in the labor market for over five years. Working with mentors with managerial experience and a longer tenure in technology also leads to high treatment effects. In the case of Challenges, participants who have just entered the labor market without a STEM degree exhibit high treatment effects. Different patterns of heterogeneity in treatment effects combined with insights from our discussions with Dare IT, mentors, and industry experts suggest that the two programs support participants’ employability through different mechanisms. Mentoring can positively impact the chances of obtaining employment in the technology sector by reinforcing self-efficacy, improving interview skills, or expanding professional networks. In contrast, creating portfolio items in the Challenges program that constitute signals of specific skills relevant to the job can counter a pessimistic belief (e.g., based on stereotypes) held by hiring managers, recruiters, or applicants themselves.
The core motivation for creating the Challenges program was to extend Dare IT’s impact beyond what is possible via the relatively small Mentoring program. Although Challenges expanded the number of people who receive Dare IT services, the existing capacity constraints in Mentoring and Challenges present a significant hurdle, as the number of interested applicants greatly exceeds the number of available spots in each program. Furthermore, at the time this research was conducted, Dare IT was not using paid advertising to promote its services, so the applicants to Dare IT heard about the program through word-of-mouth recommendations or through an internet search, which underscores that the potential demand for Dare IT services is even higher. In the case of Mentoring, the constraint is caused by the limited number of mentors and the operational difficulties of scaling up the program. For the Challenges program, the number of spots has been intentionally limited to 300 to maintain appropriate participant support. [^4] The first 100 applicants were automatically granted a spot in Challenges and were removed from the analysis conducted in this paper. Even though there are plans to incrementally increase the number of available spots in future Challenges cohorts, Dare IT will continue to grapple with the decisions of which applicants to admit to the Challenges program and how to select participants for the Mentoring program in the foreseeable future.
In the presence of capacity constraints and substantial heterogeneity in treatment effects, admission rules prioritizing applicants with the highest benefits can potentially enhance Dare IT’s impact. Adapting the methodology of Athey and Wager (2021) and Zhou et al. (2023), we estimate the value of counterfactual assignment policies that maximize outcomes across all applicants, given the constraints on the number of spots available in each program. These policies are then evaluated in the test set using doubly-robust estimates obtained through cross-fitting.
We estimate that the existing policy, which assigns 15% of all applicants to Challenges and 13% in Mentoring, leads to the weighted average treatment effect of 12 percentage points increase in the probability of getting a tech job. Considering the case in which the available number of spots in each program does not change, but we change the assignment policy to prioritizes applicants with the highest predicted benefits from each program (we refer to this policy as optimal policy), we show that the average treatment effect increases to 22 percentage points. Thus, we find a 10 percentage points incremental benefit of targeting. Second, we evaluate the optimal policy at various capacity levels. While expanding capacity increases the average outcomes as applicants who previously did not participate are admitted, the incremental value of targeting over random admissions might decrease, as applicants with the highest predicted benefits are allocated first, and the marginal participants have lower predicted benefits. However, we estimate that targeting has substantial value even at high capacity levels. We show that when the Challenges program is expanded to 50% of all applicants, the incremental benefit of targeting is 5 percentage points, while when there are no capacity limits, the incremental benefit of targeting is to increase the probability of finding a job by 3 percentage points. The comparison of random and targeted admissions at the full capacity level, underscores that some participants have higher benefits from one program than from the other; thus, the value of targeting is not only in targeting between programs and out of programs, but also across the two offerings.
Dare IT ensures high-quality matching by asking mentors to choose mentees from a list of pre-screened candidates. When comparing the outcomes of applicants to Mentoring, who were selected by mentors but randomized to the control group, with those not selected, we find that mentors tend to pick promising candidates. These candidates will likely find jobs in the technology sector, even without participating in the program. However, this approach to selecting participants might not be maximizing the potential impact of Mentoring. We use the off-policy approach to evaluate an alternative policy where mentors select candidates with the highest treatment effects. We find that the policy of choosing applicants based on their treatment effects leads to an average probability of getting a job in the technology sector of 23% across all applicants to Mentoring (we exclude applicants to Challenges from this analysis), compared to 20% when selecting the most promising candidates. This finding suggests that focusing on those most likely to benefit from the program rather than those who seem most promising overall could increase the program’s effectiveness.
Finally, a non-parametric assignment policy which requires solving a discrete optimization program, might not be realistic in the context of Dare IT. Therefore, to create viable alternatives which lead to similar benefits as optimal policies but are easier to implement, we propose simple rules that define priority groups based on observed characteristics (e.g., applicants from smaller cities are prioritized for Mentoring). We then evaluate these rules using predicted treatment effects. We find that prioritization can substantially increase the two programs’ effectiveness.
@@ 2 Literature review
This paper relates to several strands of literature. First, the paper contributes to the literature on the effectiveness of mentoring in improving career outcomes. Several influential papers focus on the effectiveness of mentoring in the academic context: Dennehy and Dasgupta (2017) employ a randomized experiment to show that assigning a female mentor to a female engineering student increases the sense of belonging in the discipline, self-efficacy, motivation, and retention for women majoring in engineering in college. Resnjanskij et al. (2021), also in a randomized experiment, document a mentoring program’s positive impact on disadvantaged high-school youth’s math grades and career plans. Several papers focus on the impact of mentoring on the careers of academic faculty members. Gardiner et al. (2007) provide empirical evidence that mentored individuals are more likely to stay at a university, receive more grant income, and receive promotions more frequently. In more recent studies, Ginther et al. (2020) and Blau et al. (2010) analyze the effectiveness of a mentorship workshop to support women in research careers. In a randomized access design similar to ours, they show that individuals in the treatment group are more likely to receive tenure at their institution, have more publications and citations, and receive more federal grants. We contribute to this literature by providing evidence from a randomized trial of the efficacy of mentoring outside of the context of education or academic jobs. [^5] The impact of mentoring on career outcomes has also been studied using surveys and other techniques based on observational data. For example, Ragins and Cotton (1999), in a survey of over 600 individuals, find that mentored respondents have better career outcomes than non-mentored respondents. They also find that the gender composition of the mentoring pair influences the effectiveness of mentoring. Similarly, Aryee et al. (1996) survey a group of workers in Hong Kong and found that mentored respondents reported a higher number of promotions and were more satisfied with their work than those not mentored. Kammeyer-Mueller and Judge (2008) carry out a meta-analysis of this literature and concludes that mentoring improves job and career satisfaction.
Second, there is rich evidence evaluating the efficacy of offline labor market training programs. Card et al. (2018) carry out a comprehensive meta-analysis of the estimates from over 200 recent studies. They find that on average, the analyzed training programs have close to zero short-term effects on employment and some small positive impact in 2-3 years after completion. Programs evaluated in Card et al. (2018) generally aim at unemployed individuals looking for blue-collar jobs. [^6] Other influential papers evaluating labor training programs have a similar focus: Barnow (1987); Bloom et al. (1997); Heckman et al. (1999). In this category, programs focused on the development of skills on-the-job tend to be more effective (Sianesi, 2008; Lechner and Gerfin, 2000). On-the-job training programs do not necessarily need to take the form of long apprenticeships to be effective: Biewen et al. (2014) compare a long apprenticeship with short-term on-the-job training, combined with a job search workshop, and find that the short-term program is equally effective or more effective than the much longer program. Fein and Hamadyk (2018) use a randomized experiment to evaluate YearUp, a sectoral training program, and find a 40% increase in quarterly earnings in the medium term. We contribute to this literature by bringing evidence from a randomized trial of a training program that happens online, is focused on individuals joining a highly paid technology sector, and is very cheap to operate. Our program has on-the-job training elements as participants develop content designed by domain experts to closely resemble the content they would produce on the job.
Third, a few studies have analyzed the efficacy of Massive Open Online Courses (MOOCs) in supporting learners’ job search. For instance, Zhenghao et al. (2015) shows that a significant majority of Coursera participants reported career benefits, particularly those from economically- and academically-disadvantaged backgrounds. However, through a difference-in-differences approach, Castaño-Muñoz and Rodrigues (2021) finds no effect on wages or employment among MOOC learners in Spain and a moderate impact on job retention. Hadavand et al. (2018) compare participants in a Data Science MOOC that were just above the threshold to graduate with those just below and argue that graduation from this MOOC leads to a 30% increase in the probability of finding a job. Our study contributes to this literature by using a randomized experiment approach. We also compare an online training program with an in-person Mentoring, identifying diverse patterns of heterogeneity that suggest different priority groups for online versus offline learning.
Fourth, we contribute to the literature on skill signaling (Spence, 1978). Previous research showed that credentials demonstrating capability can significantly boost earnings Tyler et al. (2000); Hadavand et al. (2018). Furthermore, in situations where companies might be reticent about hiring less experienced candidates, establishing trustworthy indicators of ability can bring advantages at the market level Pallais (2014). Our research, through a randomized trial, sheds light on the efficacy of a program aimed at developing labor market signals. Our subjects are women seeking tech roles who can utilize skill signaling to counter possible negative stereotypes. However, since the programs we evaluate incorporate both learning and signaling, we cannot separate the signaling effect.
Fifth, we document a high degree of heterogeneity in treatment effects from both programs, which we use to define treatment assignment policies. We show that targeting based on treatment effects leads to substantially higher average outcomes than random targeting, or targeting based on lowest outcomes in the absence of treatment, or highest outcomes with treatment. In doing so, we contribute to a small but growing literature on targeting treatment based on estimated treatment effects. An early contribution to this area was made by Ascarza (2018), who using randomized experiments, documents that treatment effect-based targeting is more effective in customer retention than churn-based prediction. Similar findings have been noted in marketing and development economics and medical contexts (Hitsch and Misra (2018), Devriendt et al. (2021), Haushofer et al. (2022), Inoue et al. (2023)). Lastly, a comparison of assignment policies for financial aid application nudges by Athey et al. (2023) highlights the superiority of maximizing treatment effects over targeting based on the predicted risk of non-compliance.

### v4

@@ Abstract
Women remain underrepresented in technology despite many support programs, partly due to limited evidence on which interventions work and how to allocate scarce program slots. We evaluate two approaches to facilitating technology-sector transitions for women in Poland: Challenges, a scalable and cheap-to-operate (\$15/person) online program developing job-relevant portfolios (designed and implemented during this research), and one-on-one mentoring. The programs address distinct labor market barriers: expanding professional networks and teaching their effective use (Mentoring) versus providing credible skill signals (Challenges). Randomizing oversubscribed admissions, we find both programs substantially increase technology employment twelve months post-completion, Mentoring by 15 percentage points (p.p.) and Challenges by 11 p.p. Exploiting treatment effect heterogeneity, we show that algorithmic targeting based on predicted treatment effects achieves outcomes 86% larger than random assignment at current capacity levels. While mentors successfully identify high-benefit candidates (outcomes 48% above random selection), algorithmic allocation using observable characteristics yields an additional 11% improvement. The results demonstrate that (i) relatively low-cost interventions can meaningfully address gender gaps in technology, and (ii) data-driven allocation rules can substantially enhance program effectiveness even when expert judgment performs well. JEL: J24, J08, J16, C93
@@ 1 Introduction
An important concern for labor policymakers is identifying and implementing effective programs to support workers transitioning into high-growth sectors. The technology sector exemplifies this challenge: despite rapid employment growth and attractive wages, women hold only 30% of technology jobs worldwide.111unitednations2021 estimates the share of women in the technology sector at approximately 30%. Many programs exist to support women in technology, including NGO initiatives (Women in Technology International, Girls Who Code), corporate programs (Women in Tech from Salesforce, The Women at Microsoft), and government efforts (UN’s Girls in ICT Day, European Commission’s Women in Digital Scoreboard). While women face multiple barriers entering technology, this paper examines two specific labor market frictions. [^2] The gender gap in technology has been attributed to discrimination in hiring and promotion (reuben2014gender; moss-racusin2012science), work-family conflict and child penalties (goldin2014grand; kleven2019children; bertrand2010dynamics), and underrepresentation in STEM education pipelines (bayer2016diversity; carrell2010sex). We focus on network and information barriers because short-term labor market interventions can directly address these constraints, unlike structural factors requiring longer-term policy responses. First, women have weaker professional networks than men, limiting access to job opportunities (greguletz2019women; forret2004networking). Second, gender stereotypes lead employers to discount women’s technical abilities, creating information frictions that credible skill signals could reduce (cheryan2013stereotypical; del2022more; murciano2022missing). Understanding which constraints bind, and for whom, is critical for designing effective policy. If network barriers dominate, interventions should expand professional connections; if information frictions dominate, scalable credentialing programs may suffice. Alternatively, both frictions may bind simultaneously for different populations, necessitating multiple program types and optimal assignment rules. We provide experimental evidence distinguishing between these possibilities.
We evaluate two interventions implemented in Poland by Dare IT, a social impact organization, each targeting one of these frictions. Mentoring pairs participants with senior female technology professionals for four months of one-on-one sessions. While sessions are personalized, mentors consistently emphasize networking: facilitating introductions to hiring managers, expanding mentees’ professional connections, and teaching effective relationship leverage. Challenges is a scalable online program (\$15 per participant) we designed during this research to address skill signaling barriers. The program helps participants build portfolios demonstrating technical proficiency through six bi-weekly industry-designed assignments, culminating in complete projects, mobile app designs or React applications, that participants reference in job applications. [^3] Program development began with structured interviews with hiring managers at twelve Polish technology firms, who uniformly cited the absence of credible practical experience signals in female applicants’ profiles. We designed Challenges to be delivered entirely online, centered on peer interaction rather than scarce mentors, and based on realistic business cases to ensure employer credibility. The program has been replicated across multiple specializations (software testing, cloud computing, UI design) and delivered in Ukrainian for refugees. We exploit both programs’ substantial oversubscription to randomize access and measure employment outcomes via LinkedIn profiles over twelve months. [^4] Technology jobs include positions in technology firms (excluding finance, legal, HR, accounting roles) or technical positions elsewhere (software development, IT support, data analytics). UX and front-end programming roles are classified as technology jobs regardless of firm type.
Both programs generate large employment effects, but with markedly different patterns of treatment effect heterogeneity that reveal which constraints bind for different populations. Mentoring increases technology employment by 15 percentage points (SE 5.7, from a 46% control baseline); Challenges by 11 percentage points (SE 4.8, from 27%). Evidence supports distinct mechanisms. Mentoring increases LinkedIn connections by 52.7 (SE 20.8), a 23% increase, while Challenges shows no detectable network expansion.
The programs benefit different populations, with heterogeneity patterns aligning with each mechanism. Professional experience reveals the starkest contrast: Challenges generates 28 percentage points larger effects for participants with under five years of experience compared to those with longer careers, while Mentoring shows the opposite pattern with 20 percentage points larger effects for experienced workers. This indicates that early-career women benefit from portfolio signals that compensate for limited work history, while experienced women benefit from network expansion that leverages their accumulated professional standing. Geography provides additional evidence: Mentoring generates 26 percentage point effects for women in small towns compared to only 3 percentage points for Warsaw residents (difference: 23 percentage points, SE 14), while Challenges shows no significant geographic heterogeneity (difference: -2 percentage points, SE 12). This pattern is consistent with network barriers binding tightly outside major urban centers where local technology ecosystems are sparse, while online portfolio signals function equally regardless of location. Mentor characteristics also predict treatment effects: participants matched to mentors with managerial experience or over five years in technology show larger treatment effects, consistent with senior mentors possessing richer networks and better access to hiring decision-makers. These differential patterns demonstrate that both network barriers and information frictions constrain technology employment, but for distinct populations defined by career stage or geography.
The differential treatment effects create an allocation problem when program capacity is limited. Both programs remain substantially oversubscribed (5:1 applicant-to-spot ratios for Mentoring; 300 spots filled in four hours for Challenges), requiring selection among applicants. We evaluate counterfactual assignment policies using the methods of athey2021policy and zhou2023offline, estimating which applicants should receive which program to maximize aggregate outcomes. A targeted policy assigning 15% of applicants to each program based on predicted treatment effects achieves outcomes 86% larger than random assignment. This gain reflects two sources of value: selecting individuals with highest treatment effects and matching individuals to their optimal program type. The matching component generates persistent value: even when capacity expands to 50% per program (universal coverage), targeted allocation maintains outcomes 23% larger than random assignment. This demonstrates that complementarity between individual characteristics and program mechanisms creates allocation gains beyond simple capacity expansion, treatment effect heterogeneity enables welfare-improving assignment even when most applicants can be served.
We further compare algorithmic targeting to expert human judgment. Mentoring relies on experienced female technology professionals to select mentees from pre-screened applicants based on pre-recorded short videos and the assessment of the application materials (cover letter, resume, and registration survey). This expert selection substantially outperforms random allocation, achieving outcomes 48% larger. Yet algorithmic allocation using only six observable characteristics (education, location, age, professional experience, degree level, domain preference) achieves outcomes 63% larger than random assignment, an additional 11% improvement over mentor judgment. The two approaches, by experts and by the algorithm, achieve high-degree of concordance (70%): the algorithm largely validates expert choices while identifying additional high-benefit candidates. Despite mentors’ rich soft information, systematic patterns in treatment response remain that simple demographic targeting can exploit. The result extends recent work showing algorithms can complement expert decisions in domains from bail (kleinberg2018human) to medical diagnosis (mullainathan2022diagnosing), demonstrating that even sophisticated expert judgment leaves room for data-driven improvement when treatment effects vary systematically with observable characteristics.
The remainder of the paper proceeds as follows. Section 2 reviews related literature. Section 3 describes the programs, theoretical framework, and testable predictions. Section 4 details the experimental design and data. Section 5 presents treatment effect estimates, mechanisms, and heterogeneity analysis. Section 6 evaluates counterfactual allocation policies. Section 7 concludes.
@@ 2 Literature Review
This paper contributes to the literature on treatment effect heterogeneity and optimal policy assignment. Building on recent advances in policy learning using machine learning methods (athey2017econometrics; kitagawa2018should; athey2021policy), we examine whether data-driven targeting can improve program allocation decisions in active labor market policy. Theoretical work demonstrates that when treatment effects vary across individuals and program capacity is limited, policies that assign treatment based on predicted treatment effects can substantially improve welfare relative to random allocation or allocation based on baseline characteristics (kitagawa2018should; bhattacharya2012inferring). Empirical applications of these methods have demonstrated value across diverse domains: customer retention (ascarza2018retention), marketing (hitsch2018heterogeneous), development programs (haushofer2022targeting), agriculture (kasy2021adaptive), and medical treatment (inoue2023machine).
We extend this literature to active labor market policy. Optimal allocation based on predicted treatment effects achieves outcomes 86% larger than random assignment at 15% capacity and 23% larger even at universal coverage. Crucially, targeting based on predicted treatment effects (CATE) dominates targeting based on baseline outcomes or predicted final outcomes, demonstrating that treatment response heterogeneity does not align with baseline heterogeneity. The persistent gains at near-universal coverage reflect complementarity between individual characteristics and program type, creating allocation value beyond capacity expansion.
We further contribute to the growing literature examining whether algorithmic decision-making can complement or improve upon expert human judgment. Recent work demonstrates that machine learning algorithms can match or exceed expert performance in domains from bail decisions (kleinberg2018human) to medical diagnosis (mullainathan2022diagnosing), often by identifying systematic patterns that experts miss or cannot consistently apply. Similar to findings in online labor markets where revealing information about worker quality improves matching outcomes (pallais2014inefficient), we show that observable characteristics predict treatment effects systematically.
We demonstrate that data-driven targeting can improve upon expert human judgment. Mentor-based selection achieves treatment effects 48% larger than random allocation. Yet algorithmic targeting using only six observable characteristics achieves an additional 11% improvement. The 70% concordance between methods indicates complementarity: algorithms largely validate expert choices while identifying additional high-benefit candidates, demonstrating that systematic patterns in treatment response exist even after expert selection captures most available heterogeneity.
This paper contributes to the literature on mentoring effectiveness in improving career outcomes. Several influential papers focus on mentoring in academic contexts: dennehy2017female; resnjanskij2021can; gardiner2007show; ginther2020can; blau2010can show that mentorship increases sense of belonging among female engineering students, youth’ learning outcomes, and tenure rates, publications, and federal grants for women in research careers. [^5] The impact of mentoring on career outcomes has also been studied using observational data. ragins1999mentor survey over 600 individuals and find that mentored respondents have better career outcomes than non-mentored respondents, with effects varying by the gender composition of the mentoring pair. kammeyer2008quantitative conduct a meta-analysis concluding that mentoring improves job and career satisfaction. Mentoring programs may operate partly through expanding professional networks, which play a critical role in labor market outcomes (granovetter1973strength; ioannides2004job; pallais2016referential). We provide randomized trial evidence on mentoring effectiveness in labor market transitions outside academic settings, demonstrating large effects on technology employment for women career-changers. Crucially, we provide experimental evidence that mentoring causally expands professional networks: Mentoring increases participants’ LinkedIn connections by 52.7 (SE 20.8), a 23% increase from baseline, while the non-network-based Challenges program shows no statistically significant effect. We further examine the evaluative selection mechanism inherent in mentor-based programs, where mentors exercise discretion in choosing mentees, and demonstrate that this selection generates substantial treatment effect heterogeneity with observable characteristics of both mentors and mentees predicting treatment effects systematically.
We contribute to the extensive literature evaluating active labor market training programs. card2018works conduct a comprehensive meta-analysis of over 200 studies, finding that traditional training programs aimed at unemployed workers exhibit near-zero short-term employment effects, with modest positive impacts emerging only 2-3 years post-completion. [^6] See also heckman1999economics, card2010active, and crepon2016active for reviews of ALMP effectiveness. Programs incorporating on-the-job skill development tend to outperform classroom-based alternatives (sianesi2008differential; lechner2000microeconometric), though even intensive apprenticeships show mixed results (biewen2014effectiveness). fein2018bridging show that YearUp, a sectoral training program targeting disadvantaged youth, generates 40% earnings gains. However, recent evidence suggests that programs targeting skill signaling and information frictions generate substantially larger employment gains than traditional classroom training (abebe2021anonymity; bassi2022screening; carranza2020job; athey2024moocs). Building on the theoretical work of spence1978job, this literature demonstrates that credentials demonstrating capability can significantly boost earnings (tyler2000estimating; hadavand2018can) and generate market-level benefits when employers face uncertainty about worker quality (pallais2014inefficient). Recent evidence further suggests that training effectiveness may depend on complementary interventions: batista2022closing find that business training for female microentrepreneurs generates significant profit increases only when combined with mobile money access, highlighting that releasing multiple constraints simultaneously may be necessary for training programs to translate into economic gains.
We provide experimental evidence on training program effectiveness in technology sector transitions, a context involving higher-skilled, career-changing women rather than the unemployed workers typically studied. We document substantial short-term effects (visible within twelve months) for both programs, aligning with the signaling literature rather than the null findings in card2018works’ meta-analysis. We directly compare two program modalities: personalized one-on-one mentoring versus scalable online project-based training. This head-to-head comparison reveals complementarity rather than substitutability: the programs serve different populations and generate heterogeneous treatment effects that enable targeted assignment. Our subjects are women seeking technology roles who can utilize skill signaling to counter gender-based statistical discrimination.